OpenStudy (maxdrizner):

A box of cereal measures 3 by 8 by 10 inches. The manufacturer plans to market a smaller box that measures 2 by 7 by 8 inches. By how much will the volume be reduced?

OpenStudy (maxdrizner):

85 in3 98 in3 112 in3 128 in3

OpenStudy (campbell_st):

well find the volume of the larger box then find the volume of the smaller box and then Volume Reduction = larger box - smaller box hope it helps

OpenStudy (paki):

Calculate the old volume... v1=3×8×10= 240 Calculate the new volume now... v2=2×7×8= 112 The change in volume is the difference of the two volumes you calculated, Δv=v1−v2 = 240-112 = 128... That will tell us how much the volume was reduced by...

OpenStudy (maxdrizner):

thank you

OpenStudy (paki):

pleasure